In microbiology, dilutions are critical for getting a countable number of microbes to grow on a petri dish. If you suspect your culture of bacteria has 100 x 106 cells per mL, what would you want the final dilution to be in order to end up with 100 colonies of bacteria on a petri dish? Express your answer as an exponent rounded to two decimal places. Eg. 10.01e-3, not 0.00101. Treat colonies of bacteria (cells / mL) the way you would treat concentration (Moles / L) The formula M1V1 = M2V2 can still be used. The final concentration is the number of colonies on the petri dish. Solve for V2.
Answer #1

CFU/ml = number of colonies / dilution factor * volume plated.

Given that CFU/ml =100 x 106 cells per mL

Number of colonies = 100

Suppose 1 ml of culture was plated.

100 x 106= 100 / Dilution factor

Dilution factor = 100/100*106= 10-6

So we need to dilute the culture by 10-6 times to get 100 colonies on the agar plate.